Y=-2x+3
8x+4y=12

1.(1,1)
2.(0,0)
3.infinity
4.no solution
1
Expert's answer
2020-04-20T17:28:57-0400

8x+4y=12 => 4y = 12-8x => y= 3-2x

"\begin{cases}
 y=-2x+3\\
 y=3-2x 
\end{cases}"

y1=y2

-2x1+3 = 3-2x2

-2x1 = -2x2

x1 = x2

valid for any x => answer 3. infinity